George, Jenna and Doreen start a new company.they agree to share any profit/loss made by the company in the same ratio as their initial contributions to the start-up capital of the company.if George, Jenna and Doreen's respective contributions to the start-up capital are R250000,00 , R100000,00 , and R50000,00 and total profit in the company's first year of operation is R160000,00 how much did each of the partners receive that year?
Answer by greenestamps(13351)



The ratio of their contributions is 250:100:50 = 5:2:1. So the fractions of the total that each of them provided is

George: 5/8
Jenna: 2/8
Doreen: 1/8

Dividing the profit of R160000 into 8 equal parts makes each part R20000. So the profit for each of the three is

George: 5*R20000 = R100000
Jenna: 2*R20000 = R40000
Doreen: 1*R20000 = R20000